I'm getting "incorrect answer" for 6.99cm in part A and -6.826cm in part B. I don't know where I'm making the mistake :(

I'm getting
Answer #1

The magnification is 97 /2.20 = 44.09

but magnification is also u/v

so

u/v= 44.09

The imaging formula for the lens is

1/v + 1/u = 1/f

where u is the distance between image and lens

v is distance between lens and object

First replace u with 44.09v

1/v + 1/(41v) = 1/f

(1/v) * (1 + 1/44.09) = 1/f

1/v = 1/(1/44.09 + 1)*(1/f)

v = f*(1.0226)

Now replace v with u /44.09:

44.09/u + 1/u = 1/f

(1/u) 44.09 = 1/f


u = 44.09f
given distance between object and screen is 301 cm

u+v = 301
44.09f+f(1.0226)= 301
f= 6.672 cm

now ,

u= 1.0226* 6.672=- 6.822 cm